# Near-Optimal Channel Estimation for Dense Array Systems

### v1

@@ III Mutual Information Maximization for Channel Estimation in DASs
This section presents a MIM based method to solve the problem outlined in (3). We first establish a general framework of the proposed design. A water-filling-inspired observation-matrix-design is then investigated as an ideal case of the proposed framework.
III-A Framework of the Proposed Channel Estimation Design
The framework of our proposed channel estimator is illustrated in Fig. 3. It consists of two stages: the MIM enabled observation matrix design and the Bayesian regression based channel estimation, each of which is explained in detail below.
III-A1 Bayesian regression based channel estimation
As previously discussed, a dense array features highly correlated channels across different antenna ports, which motivates us to recover ${\mathbf{h}}$ using Bayesian regression. Suppose the channel, ${\mathbf{h}}$, is sampled from a Gaussian process $\mathcal{CN}({\mathbf{0}}_{M},\bm{\Sigma}_{\bf h})$. The covariance matrix $\bm{\Sigma}_{\bf h}\in\mathbb{C}^{M\times M}$ characterizes our prior knowledge to the kernel of channel, which is thus referred to as the prior kernel in the sequel. Given the Gaussian distributions of the channel and noise, the joint probability distribution of $[{\mathbf{h}}^{T},{\mathbf{y}}^{T}]^{T}$ is
[…]
Thereby, the posterior mean and posterior covariance of $\bf h$ can be calculated as:
${{\bm{\mu}}}_{{\mathbf{h}}|{\mathbf{y}}}={{\bf{\Sigma}}_{\bf{h}}}{{\bf{W}}}{\left({\bf W}^{H}{\bf{\Sigma}}_{\bf{h}}{\bf W}+\sigma^{2}{\bf I}_{Q}\right)^{-1}}{{\bf{y}}},$ (5)
For an arbitrary observation matrix $\bf W$, the optimal Bayesian estimation of ${\mathbf{h}}$ is thus the posterior mean ${{\bm{\mu}}}_{{\mathbf{h}}|{\mathbf{y}}}$, i.e., $\hat{{\mathbf{h}}}={{\bm{\mu}}}_{{\mathbf{h}}|{\mathbf{y}}}$. In the subsequent discussion, ${{\bf{\Sigma}}}_{{\mathbf{h}}|{\mathbf{y}}}$ is also referred to as the posterior kernel.
III-A2 MIM enabled observation matrix design
@@
It is evident that the posterior kernel ${{\bf{\Sigma}}}_{{\mathbf{h}}|{\mathbf{y}}}$ relies heavily on the choice of the observation matrix ${\mathbf{W}}$. This fact implies that customizing an observation matrix ${\mathbf{W}}$ to the kernel structure of a dense array channel can significantly reduce the estimation error, as measured by $\mathsf{E}\|{\mathbf{h}}-\hat{{\mathbf{h}}}\|_{2}^{2}={\mathsf{Tr}}({{\bf{\Sigma}}}_{{\mathbf{h}}|{\mathbf{y}}})$. Unfortunately, the direct optimization of $\mathsf{Tr}({{\bf{\Sigma}}}_{{\mathbf{h}}|{\mathbf{y}}})$ is non-convex and intractable, making it hard to gain insights from this optimization process. To overcome this challenge, we adopt the idea of Gaussian Process Regression to design ${\mathbf{W}}$ from the perspective of MIM [41, 42, 43]. Specifically, the main objective of Gaussian Process Regression is to extract as much information about the unknown channel ${\mathbf{h}}$ from the observation ${\mathbf{y}}$ as possible. Accordingly, the observation matrix, ${\mathbf{W}}$, is designed to maximize the MI between ${\mathbf{y}}$ and ${\mathbf{h}}$:
$\max_{{\mathbf{W}}\in\mathcal{W}}I({\mathbf{y}};{\mathbf{h}})$ $=H({\mathbf{y}})-H({\mathbf{y}}|{\mathbf{h}})$
$=\log\det\left({\mathbf{I}}_{M}+\frac{1}{\sigma^{2}}{\mathbf{W}}^{H}{\bf\Sigma_{h}}{\mathbf{W}}\right),$ (7)
where $I(\cdot;\cdot)$ denotes the MI; $H(\cdot)$ denotes the entropy of its argument; and $\det(\cdot)$ denotes the determinant of its argument. It is notable that the observation matrix design in (III-A2) resembles the well-know single-user MIMO precoding problem [45]. Comprehensive comparisons between these two methodologies are provided below.
  • Category of unknown data: Both MIMO precoding and observation matrix design in (III-A2) aim to reduce the uncertainty between the received data, ${\mathbf{y}}$, and the unknown data. The former methodology regards the transmitted symbol vector, denoted by ${\mathbf{x}}$, as the unknown data. To facilitate MIMO precoding, multiple RF chains are employed to support an equal number of data streams [45]. On the other hand, the unknown data to the latter methodology is the channel, ${\mathbf{h}}$. During channel estimation, especially in our considered system, a single RF chain accompanied by repeated pilot transmission in multiple timeslots is required to recover ${\mathbf{h}}$. In comparison, one can discover that the multiple RF chains in MIMO precoding are converted to the multi-timeslot pilot resources in channel estimation. Thereby, the design of the observation matrix in (III-A2) can be viewed as a time-domain duality of classic MIMO precoding.
  • Constraint on the observation matrix: The mathematical difference between MIMO precoding and observation matrix design is primarily attributed to the constraints imposed on ${\mathbf{W}}$. For MIMO precoding, each column of ${\mathbf{W}}$ refers to a precoding vector associated with one dedicated RF chain. The optimization of precoding vectors, leveraging multiple RF chains, typically enforces the total-power-constraint on ${\mathbf{W}}$, i.e., $\|{\mathbf{W}}\|_{F}^{2}=Q$. In terms of the observation matrix design, each column of ${\mathbf{W}}$ signifies an observation vector used for receiving one pilot signal. As discussed in Section II, since each observation vector ${\mathbf{w}}_{q}$ amplifies both the desired signal and the noise, it is subject to the pilot-wise power constraint, i.e., $\|{\mathbf{w}}_{q}\|_{2}^{2}=1$ or $|w_{q,m}|=\frac{1}{\sqrt{M}}$ [27]. As a result, the distinction in the observation matrix constraint differentiates the MIM in (III-A2) from the classic MIMO precoding.
To summarize, the proposed channel estimation strategy, as illustrates in Fig. 3, involves obtaining the optimal observation matrix by solving the MIM problem (III-A2), performing pilot transmissions as presented in (2), and finally recovering the channel ${\mathbf{h}}$ using the Bayesian regression (5). It is evident that the main technical challenge in our framework comes from solving the MIM problem defined in (III-A2).
III-B Water-Filling Inspired Ideal Observation Matrix Design
To help understanding of the proposed framework, we relax the pilot-wise power constraint, $\|{\mathbf{w}}_{q}\|^{2}_{2}=1$, to the total power constraint, $\|{\mathbf{W}}\|_{F}^{2}=\sum_{q=1}^{Q}\|{\mathbf{w}}_{q}\|^{2}_{2}=Q$, to derive an ideal solution to the MIM problem in (III-A2). Note that the feasible set built on $\|{\mathbf{w}}_{q}\|^{2}_{2}=1$ is a subset of that established by $\sum_{q=1}^{Q}\|{\mathbf{w}}_{q}\|^{2}_{2}=Q$. Thereby, the derived ideal solution is practically unachievable but can serve as an upper-bound for estimation accuracy evaluation. In this context, the problem (III-A2) is transformed to the well-known single-user MIMO precoding problem, which can be globally optimized using the water-filling principle [46] as summarized in Algorithm 1.
Specifically, let the eigenvalue decomposition of ${\bf\Sigma_{h}}$ be ${\mathbf{U}}_{K}{\bf\Lambda}_{K}{\mathbf{U}}_{K}^{H}$, where $K$ is the rank of $\bf\Sigma_{h}$, ${\mathbf{U}}_{K}=[{\mathbf{u}}_{1},{\mathbf{u}}_{2},\cdots,{\mathbf{u}}_{K}]$ encompasses the eigenvectors, and ${\bf\Lambda}_{K}={\rm diag}\{{\lambda}_{1},{\lambda}_{2},\cdots,{\lambda}_{K}\}$ represents the positive eigenvalues in a descending order. The water-filling algorithm dictates that the optimal observation matrix ${\mathbf{W}}$ is constructed by the eigenvectors allocated by different powers, i.e., ${\mathbf{W}}={\mathbf{U}}_{K}{\mathbf{P}}$, with ${\mathbf{P}}\in\mathbb{R}^{K\times Q}$ representing the power allocation matrix. We assume that the pilot length $Q$ is larger than the rank $K$. Then power allocation matrix is expressed as ${\mathbf{P}}=[{\mathbf{P}}_{K},{\mathbf{0}}_{K\times(Q-K)}]$ with ${\mathbf{P}}_{K}={\rm diag}\{\sqrt{p_{1}},\sqrt{p_{2}},\cdots,\sqrt{p_{K}}\}$. The power $p_{k}$ allocated to each eigenvector is calculated by the water-filling principle:
$p_{k}=\left(\beta-\frac{\sigma^{2}}{\lambda_{k}}\right)^{+},\forall k\in\{1,2,\cdots,K\},$ (8)
where the water-level $\beta$ can be efficiently determined via binary search to meet the total power constraint: $\sum_{k=1}^{K}\left(\beta-\frac{\sigma^{2}}{\lambda_{k}}\right)^{+}=Q$.
Algorithm 1 Water-Filling-Based Observation Matrix Design
0:  Number of pilots $Q$, prior kernel ${\bm{\Sigma}}_{\bf h}$.
0:  Designed observation matrix ${\bf W}$.
1:  Eigenvalue decomposition: ${\bf\Sigma_{h}}={\mathbf{U}}_{K}{\bf\Lambda}_{K}{\mathbf{U}}_{K}^{H}$
2:  Perform water-filling to get ${\mathbf{P}}$ by (8)
3:  Construct observation matrix: ${\mathbf{W}}={\mathbf{U}}_{K}{\mathbf{P}}$
4:  return  Designed observation matrix ${\bf W}$
IV Practical Observation Matrix Design
This section elaborates on the observation matrix design in practical scenarios. Firstly, a greedy-method based principle is leveraged to solve the MIM problem in (III-A2). Building on this principle, we propose an ice-filling algorithm for designing the amplitude-and-phase controllable observation matrices. Additionally, we conduct a comprehensive analysis on the relationship between ice-filling and water-filling. Subsequently, a MM-based algorithm is proposed for the design of phase-only controllable observation matrices.
IV-A Observation Matrix Design Using Greedy Method
@@
In practical pilot transmission scenarios where the pilot-wise power constraint is imposed, the MIM problem in (III-A2) is non-convex since the objective function is non-concave with respect to (w.r.t) ${\mathbf{W}}$. To overcome this obstacle, we harness a greedy method to obtain the observation vectors in a pilot-by-pilot manner. Specifically, we define ${\bf{W}}_{t}=[{\bf{w}}_{1},{\bf{w}}_{2},\cdots,{\bf{w}}_{t}]$ as the observation matrix for timeslots $1\sim t$, where $t<Q$, and denote ${\mathbf{y}}_{t}={\mathbf{W}}_{t}^{H}{\mathbf{h}}+{\mathbf{z}}_{t}$ as the corresponding received signal. Given the current observation matrix ${\bf{W}}_{t}$, our greedy method aims to determine the next observation vector, ${\mathbf{w}}_{t+1}$, by maximizing the MI increment from timeslot $t$ to $t+1$, i.e., $\max_{{\mathbf{w}}_{t+1}}I({\mathbf{y}}_{t+1};{\mathbf{h}})-I({\mathbf{y}}_{t};{\mathbf{h}})$. As proved in Appendix A, the MI increment is expressed as
@@
Thereafter, the optimal ${\bf{w}}_{t+1}$ can be attained by solving the following quadratic problem:
@@
Before delving into the detail this problem, it is notable that repeatedly calculating the inverse of ${{{\bf{W}}_{t}^{H}}{{{\bf{\Sigma}}_{\bf{h}}}{{{{{\bf{W}}_{t}}}}}+{\sigma^{2}}{{\bf{I}}_{t}}}}$ in each timeslot $t$ is necessary for obtaining the kernel ${\bf{\Sigma}}_{t}$ in (10) and (11). To avoid these computationally inefficient matrix-inversion operations, Lemma 1 is derived to offer an efficient approach to calculate ${\bf{\Sigma}}_{t}$.
@@
As a result, equations (9) to (12) inspire us to propose the framework presented in Fig. 3 for the design of observation matrix. The framework begins with setting ${\bm{\Sigma}}_{0}$ as the prior kernel of the channel, ${\bm{\Sigma}}_{\bf h}$. The first observation vector ${\mathbf{w}}_{1}$ is then obtained by solving the quadratic problem in (11). For subsequent timeslots $t$ ($t\in\{1,2,\cdots,Q-1\}$), the posterior kernels ${\bm{\Sigma}}_{t}$ are updated from ${\bm{\Sigma}}_{t-1}$ and ${\bf w}_{t}$ using Lemma 1. The update of the posterior kernels then triggers the update of the observation vectors ${\bf w}_{t+1}$ by solving problem (11) pilot-by-pilot. These two updating steps are repeated until all $Q$ observation vectors are generated for performing the ensuing Bayesian regression.
In the sequel, we will elaborate on solving problem (11) under the amplitude-and-phase controllable and phase-only controllable cases, respectively, to complete the design of observation matrix.
@@ IV-B Ice-Filling for Amplitude-and-Phase Controllable Observation Matrix
7:     Eigenvalue-preserve: $\lambda_{k}^{t+1}=\lambda_{k}^{t}$ with $k\in\{1,2,\cdots,K\}-k_{t}$
@@
For amplitude-and-phase controllable observation matrices, problem (11) can be formulated as the well-known Rayleigh quotient problem:
@@
where the principal eigenvector of ${\bf{\Sigma}}_{t}$ yields the optimal solution to ${\mathbf{w}}_{t+1}$, i.e., ${\bf{\Sigma}}_{t}{\bf{w}}_{t+1}=\lambda_{\max}({\bf{\Sigma}}_{t}){\bf{w}}_{t+1}$. By combining the updating strategy of ${\bm{\Sigma}}_{t}$ described in Lemma 1 and the property of the principal eigenvector ${\mathbf{w}}_{t+1}$, we can prove Theorem 1, which lays the foundation for our ice-filling algorithm in designing amplitude-and-phase controllable observation matrices.
@@
Motivated by Theorem 1, our proposed ice-filling algorithm is intrinsically an eigenvector-assignment-process, as presented in Algorithm 2. To elaborate, we define $\{\lambda_{1}^{t},\lambda_{2}^{t},\cdots,\lambda_{K}^{t}\}$ as the eigenvalues of the posterior kernel ${\bm{\Sigma}}_{t}$, which are initialized as the eigenvalues of the prior kernel ${\bm{\Sigma}}_{\bf h}$ in Step 2. In each timeslot, we identify the largest eigenvalue from $\{\lambda_{1}^{t},\lambda_{2}^{t},\cdots,\lambda_{K}^{t}\}$ and index it by $k_{t}$ in Step 4. Then, the corresponding eigenvector ${\mathbf{u}}_{k_{t}}$ of ${\bm{\Sigma}_{h}}$, which is equivalent to the principal eigenvector of ${\bm{\Sigma}}_{t}$, is assigned to ${\mathbf{w}}_{t+1}$ in Step 5. Finally, according to Theorem 1, the selected eigenvalue $\lambda_{k_{t}}^{t}$ is squeezed to $\frac{\lambda_{k_{t}}^{t}\sigma^{2}}{\lambda_{k_{t}}^{t}+\sigma^{2}}$, while the other eigenvalues of ${\bm{\Sigma}}_{t}$ are preserved, to attain the eigenvalues of the next kernel ${\bm{\Sigma}}_{t+1}$. These steps are repeatedly executed until all observation vectors are generated, which completes the proposed ice-filling algorithm.

### v2

@@ III Mutual Information Maximization for Channel Estimation in DASs
In this section, we first presents a MIM based framework for designing observation matrices. Then, a water-filling-inspired observation-matrix-design is investigated as an ideal case of the proposed framework.
III-A MIM based framework for observation matrix design
We adopt the MI between the received signal ${\mathbf{y}}$ and the channel ${\mathbf{h}}$ to evaluate the quality of the designed observation matrix. There are two reasons for using the MI as a metric. Firstly, MI tells the amount of information about the unknown channel ${\mathbf{h}}$ we can gain from the received signal ${\mathbf{y}}$. The second reason is that, according to the information-theoretic properties of Bayesian statistics from [46], the maximization of MI is asymptotically the minimization of Cramer Rao Bound (CRB).
Given the distributions ${\mathbf{h}}\sim\mathcal{CN}({\mathbf{0}}_{M},\bm{\Sigma}_{\bf h})$ and ${\bf z}\sim\mathcal{CN}({\mathbf{0}}_{Q},\sigma^{2}{\mathbf{I}}_{Q})$, the MI is formulated as
$\max_{{\mathbf{W}}\in\mathcal{W}}I({\mathbf{y}};{\mathbf{h}})$ $=\log\det\left({\mathbf{I}}_{Q}+\frac{1}{\sigma^{2}}{\mathbf{W}}^{H}{\bf\Sigma_{h}}{\mathbf{W}}\right),$ (6)
where $I(\cdot;\cdot)$ denotes the MI, $\det(\cdot)$ is the determinant of its argument, and $\mathcal{W}$ represents the feasible set of ${\mathbf{W}}$ for both amplitude-and-phase controllable and phase-only controllable analog combiners. It is notable that the observation matrix design problem in (6) resembles the well-know point-to-point MIMO precoding problem [47]. Comprehensive comparisons between these two methodologies are provided below.
  • Category of unknown data: Both MIMO precoding and observation matrix design in (6) aim to reduce the uncertainty between the received data, ${\mathbf{y}}$, and the unknown data. The former methodology regards the transmitted symbol vector, denoted by ${\mathbf{x}}$, as the unknown data. To facilitate MIMO precoding, multiple RF chains are employed to support an equal number of data streams. On the other hand, the unknown data to the latter methodology is the channel, ${\mathbf{h}}$. Our considered uplink channel estimation requires transmitting pilot signals over multiple timeslots to recover ${\mathbf{h}}$. In comparison, one can discover that the multiple RF chains in MIMO precoding are converted to the multi-timeslot pilot resources in channel estimation. Thereby, the design of the observation matrix in (6) can be viewed as a time-domain duality of point-to-point MIMO precoding.
  • Constraint on the observation matrix: The mathematical difference between MIMO precoding and observation matrix design for uplink channel estimation is attributed to the constraints of ${\mathbf{W}}$. For MIMO precoding, each column of ${\mathbf{W}}$ refers to a precoding vector associated with one dedicated RF chain. The optimization of precoding vectors, leveraging multiple RF chains, typically enforces the total-power-constraint on ${\mathbf{W}}$, i.e., $\|{\mathbf{W}}\|_{F}^{2}=Q$. In terms of the observation matrix design, each column of ${\mathbf{W}}$ signifies an observation vector used for receiving one pilot signal. As discussed in Section II, since each observation vector ${\mathbf{w}}_{q}$ amplifies both the desired signal and the noise, it is subject to the pilot-wise power constraint, i.e., $\|{\mathbf{w}}_{q}\|_{2}^{2}=1$ or $|w_{q,m}|=\frac{1}{\sqrt{M}}$. The distinction in the observation matrix constraint differentiates the MIM in (6) from the classical MIMO precoding.
III-B Water-Filling Inspired Ideal Observation Matrix Design
Prior to addressing problem (6), we would like to consider an ideal (but practically unachievable) situation to establish an upper bound of the channel estimation performance. Specifically, we temporarily make the ideal assumption that the noise vector ${\mathbf{z}}$ in (2) is independent of the observation matrix ${\mathbf{W}}$ and its distribution is always $\mathcal{CN}({\mathbf{0}}_{Q},\sigma^{2}{\mathbf{I}}_{Q})$ regardless of the power of ${\mathbf{w}}_{q}$. Under this assumption, the pilot-wise power constraint, $\|{\mathbf{w}}_{q}\|^{2}_{2}=1$, can be relaxed to the total power constraint, $\|{\mathbf{W}}\|_{F}^{2}=\sum_{q=1}^{Q}\|{\mathbf{w}}_{q}\|^{2}_{2}=Q$. In this context, problem (6) is identical to the point-to-point MIMO precoding problem, which can be optimally solved using the water-filling principle [48]. It is crucial to emphasize again that this ideal situation is not practical because the noise vector ${\mathbf{z}}$ is inevitably amplified by observation matrix ${\mathbf{W}}$ during uplink channel estimation. This situation is only utilized to show the performance upper bound of channel estimation.
Let the eigenvalue decomposition of ${\bf\Sigma_{h}}$ be ${\mathbf{U}}_{K}{\bf\Lambda}_{K}{\mathbf{U}}_{K}^{H}$, where $K$ is the rank of $\bf\Sigma_{h}$, ${\mathbf{U}}_{K}=[{\mathbf{u}}_{1},{\mathbf{u}}_{2},\cdots,{\mathbf{u}}_{K}]$ encompasses the eigenvectors, and ${\bf\Lambda}_{K}={\rm diag}\{{\lambda}_{1},{\lambda}_{2},\cdots,{\lambda}_{K}\}$ represents the eigenvalues in a descending order. The water-filling algorithm dictates that the optimal observation matrix ${\mathbf{W}}$ is ${\mathbf{W}}={\mathbf{U}}_{K}{\mathbf{P}}$, where ${\mathbf{P}}\in\mathbb{R}^{K\times Q}$ represents the power allocation matrix. We assume that the pilot length $Q$ is larger than the rank $K$. Then power allocation matrix is expressed as ${\mathbf{P}}=[{\mathbf{P}}_{K},{\mathbf{0}}_{K\times(Q-K)}]$ with ${\mathbf{P}}_{K}={\rm diag}\{\sqrt{p_{1}},\sqrt{p_{2}},\cdots,\sqrt{p_{K}}\}$. The power $p_{k}$ allocated to each eigenvector is calculated by the water-filling principle:
$p_{k}=(\beta-{\sigma^{2}}/{\lambda_{k}})^{+},\forall k\in\{1,2,\cdots,K\},$ (7)
where the water-level $\beta$ is adjusted to meet the total power constraint: $\sum_{k=1}^{K}p_{k}=Q$.
IV Practical Observation Matrix Design
This section elaborates on the observation matrix design in practical scenarios. Firstly, a greedy-method based principle is leveraged to solve the MIM problem in (6). Building on this principle, we propose an ice-filling algorithm for designing the amplitude-and-phase controllable observation matrices. Additionally, we conduct a comprehensive analysis on the relationship between ice-filling and water-filling. Subsequently, a MM-based algorithm is proposed for the design of phase-only controllable observation matrices. Last, the choice of prior kernel is also discussed.
IV-A Observation Matrix Design Using Greedy Method
@@
In practical uplink channel estimation where the noise vector ${\mathbf{z}}$ is amplified by the observation vector ${\mathbf{w}}_{q}$, the pilot-wise power constraint should be considered. In this case, the MIM problem in (6) is an intractable non-convex programming. To overcome this challenge, we harness a greedy method to obtain the observation vectors in a pilot-by-pilot manner, as shown in Fig. 2. Define ${\bf{W}}_{t}=[{\bf{w}}_{1},{\bf{w}}_{2},\cdots,{\bf{w}}_{t}]$ as the observation matrix for timeslots $1\sim t$, where $t<Q$, and denote ${\mathbf{y}}_{t}={\mathbf{W}}_{t}^{H}{\mathbf{h}}+{\mathbf{z}}_{t}$ as the corresponding received signal. Given the current observation matrix ${\bf{W}}_{t}$, our greedy method aims to determine the next observation vector, ${\mathbf{w}}_{t+1}$, by maximizing the MI increment from timeslot $t$ to $t+1$, i.e., $\max_{{\mathbf{w}}_{t+1}}I({\mathbf{y}}_{t+1};{\mathbf{h}})-I({\mathbf{y}}_{t};{\mathbf{h}})$. As proved in [49], the MI increment is expressed as
@@
Thereafter, the optimal ${\bf{w}}_{t+1}$ can be attained by solving the quadratic problem:
@@
Before delving into the detail this problem, it is notable that repeatedly calculating the inverse of ${{{\bf{W}}_{t}^{H}}{{{\bf{\Sigma}}_{\bf{h}}}{{{{{\bf{W}}_{t}}}}}+{\sigma^{2}}{{\bf{I}}_{t}}}}$ in each timeslot $t$ is necessary for obtaining the kernel ${\bf{\Sigma}}_{t}$ in (9) and (10). To avoid these computationally inefficient matrix-inversion operations, Lemma 1 would be helpful to offer an efficient approach to obtain ${\bf{\Sigma}}_{t}$.
@@
As a result, equations (8) to (11) inspire us to propose the framework presented in Fig. 2 for the design of observation matrix. The framework begins with setting ${\bm{\Sigma}}_{0}$ as the prior kernel of the channel, ${\bm{\Sigma}}_{\bf h}$. The first observation vector ${\mathbf{w}}_{1}$ is obtained by solving the quadratic problem in (10). For subsequent timeslots $t$ ($t\in\{1,2,\cdots,Q-1\}$), the posterior kernels ${\bm{\Sigma}}_{t}$ are updated from ${\bm{\Sigma}}_{t-1}$ and ${\bf w}_{t}$ using Lemma 1. The update of the posterior kernels then triggers the update of the observation vectors ${\bf w}_{t+1}$ by solving problem (10) pilot-by-pilot. These two updating steps are repeated until all $Q$ observation vectors are generated for performing the ensuing Bayesian regression.
In the sequel, we will elaborate on solving problem (10) under the amplitude-and-phase controllable and phase-only controllable analog combiners, respectively.
@@ IV-B Ice-Filling for Amplitude-and-Phase Controllable Observation Matrix
7:     Eigenvalue-preserve: $\lambda_{k}^{t+1}=\lambda_{k}^{t}$ with $k\in\{1,2,\cdots,K\}/\{k_{t}\}$
@@
For amplitude-and-phase controllable observation matrices, problem (10) is the well-known Rayleigh quotient problem:
@@
where the principal eigenvector of ${\bf{\Sigma}}_{t}$ yields the optimal solution to ${\mathbf{w}}_{t+1}$, i.e., ${\bf{\Sigma}}_{t}{\bf{w}}_{t+1}=\lambda_{\max}({\bf{\Sigma}}_{t}){\bf{w}}_{t+1}$. By leveraging the updating strategy of ${\bm{\Sigma}}_{t}$ described in Lemma 1 and the property of eigenvector, we can prove Theorem 1, which lays the foundation for our ice-filling algorithm.
@@
Motivated by Theorem 1, our proposed ice-filling algorithm is intrinsically an eigenvector-assignment-process, as presented in Algorithm 1. To elaborate, we define $\{\lambda_{1}^{t},\lambda_{2}^{t},\cdots,\lambda_{K}^{t}\}$ as the eigenvalues of the posterior kernel ${\bm{\Sigma}}_{t}$, which are initialized as the eigenvalues of the prior kernel ${\bm{\Sigma}}_{\bf h}$ in Step 2. In each timeslot, we find the largest eigenvalue from $\{\lambda_{1}^{t},\lambda_{2}^{t},\cdots,\lambda_{K}^{t}\}$ and index it by $k_{t}$ in Step 4. Then, the corresponding eigenvector ${\mathbf{u}}_{k_{t}}$ of ${\bm{\Sigma}_{h}}$, which is equivalent to the principal eigenvector of ${\bm{\Sigma}}_{t}$, is assigned to ${\mathbf{w}}_{t+1}$ in Step 5. Finally, according to Theorem 1, the selected eigenvalue $\lambda_{k_{t}}^{t}$ is squeezed to $\frac{\lambda_{k_{t}}^{t}\sigma^{2}}{\lambda_{k_{t}}^{t}+\sigma^{2}}$, while the other eigenvalues of ${\bm{\Sigma}}_{t}$ are preserved, to attain the eigenvalues of the next kernel ${\bm{\Sigma}}_{t+1}$. These steps are repeatedly executed until all observation vectors are generated, which completes the ice-filling algorithm.